OpenStudy (anonymous):

How to Solve arctan(-sqrt3/3)? Please help! Got a trig test tommorow :/

OpenStudy (anonymous):

\[ \tan(\theta) = -\frac{\sqrt 3}{3} \]

OpenStudy (anonymous):



OpenStudy (anonymous):